# Locally Private Sampling with Public Data

### v1

@@ Abstract
Local differential privacy (LDP) is increasingly employed in privacy-preserving machine learning to protect user data before sharing it with an untrusted aggregator. Most LDP methods assume that users possess only a single data record, which is a significant limitation since users often gather extensive datasets (e.g., images, text, time-series data) and frequently have access to public datasets. To address this limitation, we propose a locally private sampling framework that leverages both the private and public datasets of each user. Specifically, we assume each user has two distributions: $p$ and $q$ that represent their private dataset and the public dataset, respectively. The objective is to design a mechanism that generates a private sample approximating $p$ while simultaneously preserving $q$. We frame this objective as a minimax optimization problem using $f$-divergence as the utility measure. We fully characterize the minimax optimal mechanisms for general $f$-divergences provided that $p$ and $q$ are discrete distributions. Remarkably, we demonstrate that this optimal mechanism is universal across all $f$-divergences. Experiments validate the effectiveness of our minimax optimal sampler compared to the state-of-the-art locally private sampler.
@@ 1 Introduction
Differential privacy (DP) has become the de facto standard for ensuring privacy in machine learning, and it has been widely adopted by major technology companies such as Google [17, 12], Microsoft [16], LinkedIn [32], and Meta [38]. Intuitively, an algorithm is considered differentially private if small changes to the input data—such as modifying a single entry—do not significantly affect the algorithm’s output. This traditional model assumes that a trusted curator has full access to the dataset, which can be a limiting factor in many practical scenarios.
@@
To relax this assumption, local differential privacy (LDP) was introduced by Kasiviswanathan et al. [27], allowing users to randomize their data on their own devices before sharing it with an untrusted curator. This model, with considerably weaker assumption in the trust model, has made LDP particularly appealing to both users and tech companies — the first large-scale deployments of DP was in the local model [15, 17].
@@
However, a significant limitation of traditional LDP approaches is the assumption that each user only possesses a single data record. In reality, users typically have multiple data records across different modalities, such as photos, text messages, or even whole datasets. To address this issue, Husain et al. [23] proposed treating each user’s local data as a probability distribution. They then developed a differentially private sampling algorithm that takes a user’s probability distribution as input and releases a single sample that closely approximates the distribution. This method demonstrated better performance compared to ad-hoc private sampling techniques such as private kernel density estimation [4] and private GAN [36]. Nevertheless, it still has a major limitation: it relies heavily on an arbitrary reference distribution. As a result, the performance of the algorithm is highly sensitive to the choice of reference distribution, making it less reliable in practice.
In our work, we resolve this issue by identifying the optimal target distribution for the worst-case input distribution. Our approach ensures a minimal gap between the worst-case input and the corresponding sampling distribution, guaranteeing robust performance. This makes our locally private sampler more versatile, as it can be applied regardless of the underlying distribution. Furthermore, we recognize that many applications involve publicly available data that do not require privacy protection. Incorporating such public data might improve the overall accuracy of the system, as evidenced by several recent works (see Section 1.1). Since public data is not subject to privacy guarantees, our private sampler should utilize it without perturbing it.
@@
  • As a special case, we delineate that, the randomized response mechanism [35] is optimal when the public prior is the uniform distribution.
  • We demonstrate the considerably better performance of our approach through comprehensive benchmarks against the baseline method proposed by Husain et al. [23] on synthetic and real-world datasets.
@@ 1.1 Related Work
Local differential privacy (LDP) has been widely applied to various statistical problems, such as distribution estimation [7, 37, 2, 25, 8, 3, 18, 33, 19]. In all of these works, LDP mechanisms are typically applied to a single data record per user, an assumption that overlooks practical cases where users possess datasets consisting of multiple records across different modalities.
A closely related problem to private distribution estimation is the private sampling problem, which focuses on privately generating a single sample from the underlying distribution, rather than learning the distribution itself. Raskhodnikova et al. [31] studied private sampling under the total variation distance in the central model for discrete distributions, and this was later extended to multi-dimensional Gaussian distributions by Ghazi et al. [21].
In the local model, Husain et al. [23] introduced an efficient algorithm for privately sampling from a distribution $p$ by constructing a small ball of a certain radius, referred to as a “relative mollifier”, around a reference distribution $q$, and projecting $p$ onto this ball in terms of ${\mathsf{KL}}$-divergence (see Section 2.2 for further details). It is worth mentioning that the reference distribution $q$ can be interpreted as a public prior as the algorithm keeps it invariant (similar to our construction).
Flemings et al. [20] extended the relative mollifier concept to study Rényi differential privacy guarantees in the training of large language models. Their approach demonstrated superior performance compared to standard DP stochastic gradient methods [1] on large-scale datasets, highlighting the broader applicability of private sampling techniques.
Recently, there has been a growing interest in incorporating public data into private learning problems (e.g., [29, 30, 6, 10, 11, 5, 39, 28, 34, 26, 9] to name a few). While most of these studies focus on the central DP model, notable exceptions include [29, 30], which explore the role of public data in private mean estimation and regression, respectively. To the best of our knowledge, our work is the first to address the problem of private sampling with the inclusion of public data in the local model.
@@ 2.2 Mollifiers
Definition 2 (Relative Mollifiers, [23]).
@@ 2.3 Locally Private Samplers
Definition 3 (Private Samplers [23]).
@@
This definition closely mirrors the standard formulation of LDP [27], with the key difference that the input to the algorithm is a distribution, rather than a single data point. As highlighted by Husain et al. [23], one approach to ensure that a sampler is $\varepsilon$-locally private is to guarantee that the distribution of $\mathcal{A}(p)$ lies within an $\varepsilon$-mollifier for any $p\in\Delta(\mathcal{X})$. This observation motivates the following definition.
@@
It is important to note that [23] employed a fundamentally different methodology. Their approach begins by selecting an arbitrary reference distribution $q\in\Delta(\mathcal{X})$ and constructing $\mathcal{M}_{\varepsilon,q}$, the relative $\varepsilon$-mollifier around $q$. They then project $p$ onto $\mathcal{M}_{\varepsilon,q}$ (using ${\mathsf{KL}}$-divergence), and then sample from the projection. Specifically, their sampling distribution $\hat{p}$ is obtained by solving
@@ 3 Main Results: Minimax Optimal Mechanisms
${\Gamma_{f}(q,\varepsilon)}\coloneqq\inf_{\begin{subarray}{c}{\mathsf{K}}\in{\mathcal{Q}}_{\varepsilon}\\ q{\mathsf{K}}=q\end{subarray}}\sup_{p\in\Delta({\mathcal{X}})}{D_{f}}(p\|p{\mathsf{K}}).$ (4)
The mechanism that solves the above optimization is the minimax optimal mechanism with respect to the choice of $f$-divergence. In the case of ${\mathsf{TV}}$-distance, such an optimal mechanism has a practical interpretation. Let $X\sim p$ and $Z$ is the output of an $\varepsilon$-locally private sampler. Optimal mechanism attaining $\Gamma_{{\mathsf{TV}}}(q,\varepsilon)$ corresponds to the best sampler that minimizes the worst-case $\Pr(X\neq Z)$.
@@ 3.1 Optimal Utility
${\Gamma_{f}(q,\varepsilon)}=\frac{1-q_{\min}}{e^{\varepsilon}q_{\min}+1-q_{\min}}f(0)+\frac{e^{\varepsilon}q_{\min}}{e^{\varepsilon}q_{\min}+1-q_{\min}}f\left(\frac{e^{\varepsilon}q_{\min}+1-q_{\min}}{e^{\varepsilon}q_{\min}}\right).$ (5)
@@ 3.2 Optimal Mechanisms
$\sup_{p\in\Delta({\mathcal{X}})}{D_{f}}(p\|p{\mathsf{K}})=\max_{i}{D_{f}}(\delta_{i}\|\delta_{i}{\mathsf{K}})=f(0)(1-{\mathsf{K}}_{\min})+{\mathsf{K}}_{\min}f(\frac{1}{{\mathsf{K}}_{\min}}).$ (9)
@@
Figure 1: Visualization of a single recursion step in Algorithm 1. The matrix ${\mathsf{K}}_{q,\varepsilon}$ shows the relationship between query probabilities and privacy parameters, with sub-matrix $m{\mathsf{K}}_{\bar{q},\varepsilon}$ representing internal recursive structures.
@@
$\sup_{p\in\Delta({\mathcal{X}})}{D_{f}}(p\|p{\mathsf{K}}_{q,\varepsilon})=\frac{1-q_{\min}}{e^{\varepsilon}q_{\min}+1-q_{\min}}f(0)\quad+\frac{e^{\varepsilon}q_{\min}}{e^{\varepsilon}q_{\min}+1-q_{\min}}f\left(\frac{e^{\varepsilon}q_{\min}+1-q_{\min}}{e^{\varepsilon}q_{\min}}\right).$ (14)
@@
It is important to note that our algorithm achieves a time complexity of $O(n^{3})$. This is a significant improvement over the naive linear programming approach for solving (4) in the case of ${\mathsf{TV}}$-distance. The naive method has a complexity of $O\left(n^{5}\log\left(\frac{n}{\delta}\right)\right)$ [14], where $n$ is the support size of $p$ and $q$, and $\delta$ is the desired accuracy. With an $O(n^{3})$ complexity, our method scales efficiently to much larger datasets.
@@
Let ${\mathcal{X}}$ be a finite set and let $q_{\text{u}}$ represent the uniform distribution on ${\mathcal{X}}$. Then, the $n$-ary randomized response mechanism, ${\mathsf{K}}_{\text{RR}}$ [35], is optimal in the sense:
@@ 4 Overview of Proofs
${\Gamma_{f}(q,\varepsilon)}\geq\frac{1-q_{\min}}{e^{\varepsilon}q_{\min}+1-q_{\min}}f(0)+\frac{e^{\varepsilon}q_{\min}}{e^{\varepsilon}q_{\min}+1-q_{\min}}f(\frac{e^{\varepsilon}q_{\min}+1-q_{\min}}{e^{\varepsilon}q_{\min}}).$ (16)
@@
${\mathsf{K}}_{\min}f(\frac{1}{{\mathsf{K}}_{\min}})+f(0)(1-{\mathsf{K}}_{\min})\geq\frac{1-q_{\min}}{e^{\varepsilon}q_{\min}+1-q_{\min}}f(0)+\frac{e^{\varepsilon}q_{\min}}{e^{\varepsilon}q_{\min}+1-q_{\min}}f(\frac{e^{\varepsilon}q_{\min}+1-q_{\min}}{e^{\varepsilon}q_{\min}}).$ (18)
@@
$\begin{split}\sup_{p\in\Delta({\mathcal{X}})}{D_{f}}(p\|p{\mathsf{K}}_{q,\varepsilon})=\frac{1-q_{\min}}{e^{\varepsilon}q_{\min}+1-q_{\min}}f(0)+\frac{e^{\varepsilon}q_{\min}}{e^{\varepsilon}q_{\min}+1-q_{\min}}f(\frac{e^{\varepsilon}q_{\min}+1-q_{\min}}{e^{\varepsilon}q_{\min}}).\end{split}$ (19)
@@ 5 Experiments
In this section, we compare the performance of our minimax optimal private sampler to the state-of-the-art private sampler, the relative mollifier framework (Definition 2). The evaluation is done on synthetic data and two real-world datasets: Avezu CTR [24] and MovieLens 1M [22]. Details on the computing resources are given in Section A.10.
@@ 5.1 Computation of Samplers
As for the relative mollifier sampling framework (Definition 2), we interpret the reference distribution in Definition 2 as the public prior. In their framework, we determine the normalizing constant $C$ described in (3) using a bisection approach [13]. While Husain et al. [23] focus exclusively on ${\mathsf{KL}}$-divergence in the relative mollifier framework, we extend their work to include the ${\mathsf{TV}}$-distance, offering an additional benchmark. This extension uses linear programming, with details provided in Appendix A.8.

### v2

@@ Abstract
Local differential privacy (LDP) is increasingly employed in privacy-preserving machine learning to protect user data before sharing it with an untrusted aggregator. Most LDP methods assume that users possess only a single data record, which is a significant limitation since users often gather extensive datasets (e.g., images, text, time-series data) and frequently have access to public datasets. To address this limitation, we propose a locally private sampling framework that leverages both the private and public datasets of each user. Specifically, we assume each user has two distributions: $p$ and $q$ that represent their private dataset and the public dataset, respectively. The objective is to design a mechanism that generates a private sample approximating $p$ while simultaneously preserving $q$. We frame this objective as a minimax optimization problem using $f$-divergence as the utility measure. We fully characterize the minimax optimal mechanisms for general $f$-divergences provided that $p$ and $q$ are discrete distributions. Remarkably, we demonstrate that this optimal mechanism is universal across all $f$-divergences. Experiments validate the effectiveness of our minimax optimal sampler compared to the state-of-the-art private sampler.
@@ 1 Introduction
Differential privacy (DP) has become the de facto standard for ensuring privacy in machine learning, and it has been widely adopted by major technology companies such as Google (Erlingsson et al., 2014; Bittau et al., 2017), Microsoft (Ding et al., 2017), LinkedIn (Rogers et al., 2020), and Meta (Yousefpour et al., 2022). Intuitively, an algorithm is considered differentially private if small changes to the input data—such as modifying a single entry—do not significantly affect the algorithm’s output. This traditional model assumes that a trusted curator has full access to the dataset, which can be a limiting factor in many practical scenarios.
@@
To relax this assumption, local differential privacy (LDP) was introduced by Kasiviswanathan et al. (2011), allowing users to randomize their data on their own devices before sharing it with an untrusted curator. This model, with considerably weaker assumption in the trust model, has made LDP particularly appealing to both users and tech companies — the first large-scale deployments of DP was in the local model (Differential Privacy team Apple, 2017; Erlingsson et al., 2014).
@@
However, a significant limitation of traditional LDP approaches is the assumption that each user only possesses a single data record. In reality, users typically have multiple data records across different modalities, such as photos, text messages, or even whole datasets. A line of research, called user-level privacy, addresses this by assuming that the users possess datasets of the same size generated from an underlying distribution (Levy et al., 2021; Ghazi et al., 2021; Zhao et al., 2025; Kent et al., 2024). Although this line of work is more realistic, the assumption that all users must have a dataset of the same size introduces practical limitations. To mitigate this shortcoming, Husain et al. (2020) proposed treating each user’s local data as a probability distribution. They then developed a differentially private sampling algorithm that takes a user’s probability distribution as input and releases a single sample that closely approximates the distribution. This method demonstrated better performance compared to ad-hoc private sampling techniques such as private kernel density estimation (Aldà and Rubinstein, 2017) and private GAN (Xie et al., 2018). Nevertheless, it still has a major limitation: it relies heavily on an arbitrary reference distribution. As a result, the performance of the algorithm is highly sensitive to the choice of reference distribution, making it less reliable in practice. To address these limitations, (Park et al., 2024) recently developed a minimax formulation for locally private sampling. They fully characterized the minimax risks in terms of all general $f$-divergences in both discrete and continuous settings and identified the optimal samplers that achieve these risks. Their empirical results demonstrated that their samplers significantly outperform the approach of Husain et al. (2020) for any reference distribution.
Building on this framework, we extend the work of Park et al. (2024) to incorporate publicly available data that do not require privacy protection. Our goal is to design optimal samplers that generate private samples while preserving the integrity of public data. This ensures that when the user’s data distribution closely aligns with the public data (viewed as a distribution), the resulting sampling distribution remains similarly aligned.
@@
  • As a special case, we delineate that, the randomized response mechanism (Warner, 1965) is optimal when the public prior is the uniform distribution.
  • We demonstrate the considerably better performance of our approach through comprehensive benchmarks against the baseline method proposed by Husain et al. (2020) on synthetic and real-world datasets.
@@ 1.1 Related Work
Local differential privacy (LDP) has been widely applied to various statistical problems, such as distribution estimation (Asoodeh and Zhang, 2024; Ye and Barg, 2018; Acharya et al., 2019; Kairouz et al., 2016; Barnes et al., 2020; Acharya et al., 2024; Feldman and Talwar, 2021; Shah et al., 2021; Feldman et al., 2022). In all of these works, LDP mechanisms are typically applied to a single data record per user, an assumption that overlooks practical cases where users possess datasets consisting of multiple records across different modalities.
A closely related problem to private distribution estimation is the private sampling problem, which focuses on privately generating a single sample from the underlying distribution, rather than learning the distribution itself. Raskhodnikova et al. (2021) studied private sampling under the total variation distance in the central model for discrete distributions, and this was later extended to multi-dimensional Gaussian distributions by Ghazi et al. (2024).
In the local model, Husain et al. (2020) proposed an efficient algorithm for privately sampling from a distribution $p$. Their approach involves constructing a small ball of a certain radius (referred to as a “relative mollifier”), around a reference distribution $q$ and projecting $p$ onto this ball with respect to the ${\mathsf{KL}}$-divergence (see Section 2.2 for further details). Notably, the reference distribution $q$ can be interpreted as a public prior, as the algorithm preserves its invariance, a property that aligns with our approach.
Flemings et al. (2024) extended the relative mollifier concept to study Rényi differential privacy guarantees in the training of large language models. Their approach demonstrated superior performance compared to standard DP stochastic gradient methods (Abadi et al., 2016) on large-scale datasets, highlighting the broader applicability of private sampling techniques.
More recently, Park et al. (2024) introduced a minimax formulation to determine optimal samplers under the assumption that $p$ is absolutely continuous with respect to a positive measure with its Radon-Nikodym derivative satisfying a certain regularity condition. They identified two families of optimal samplers: linear and non-linear. A linear sampler generates samples from a distribution obtained by applying an $\varepsilon$-LDP Markov kernel to $p$, while a non-linear sampler constructs the sampling distribution by projecting $p$ onto a convex set specified by $\varepsilon$.
Recently, there has been a growing interest in incorporating public data into private learning problems (e.g., Lowy et al. (2024); Ma et al. (2024); Amid et al. (2022); Ben-David et al. (2023); Bie et al. (2022); Alon et al. (2019); Yu et al. (2021); Liu et al. (2023); Wang and Zhou (2020); Kairouz et al. (2021); Bassily et al. (2020) to name a few). Although most of these studies focus on the central DP model, notable exceptions include Lowy et al. (2024); Ma et al. (2024), which explore the role of public data in private mean estimation and regression, respectively.
Motivated by these developments, we extend the framework of Park et al. (2024) in the linear setting to incorporate a public prior and identify optimal linear private samplers. Consequently, our objective reduces to determining the optimal $\varepsilon$-LDP Markov kernel ${\mathsf{K}}$ that preserves the public prior. To the best of our knowledge, our work is the first to address the problem of private sampling with the inclusion of public data in the local model.
@@ 2.2 Mollifiers
Definition 2 (Relative Mollifiers, Husain et al. (2020)).
@@ 2.3 Locally Private Samplers
Definition 3 (Private Samplers Husain et al. (2020)).
@@
This definition closely mirrors the standard formulation of LDP Kasiviswanathan et al. (2011), with the key difference that the input to the algorithm is a distribution, rather than a single data point. As highlighted by Husain et al. (2020), one approach to ensure that a sampler is $\varepsilon$-locally private is to guarantee that the distribution of $\mathcal{A}(p)$ lies within an $\varepsilon$-mollifier for any $p\in\Delta(\mathcal{X})$. This observation motivates the following definition.
@@
It is important to note that Husain et al. (2020) employed a fundamentally different methodology. Their approach begins by selecting an arbitrary reference distribution $q\in\Delta(\mathcal{X})$ and constructing $\mathcal{M}_{\varepsilon,q}$, the relative $\varepsilon$-mollifier around $q$. They then project $p$ onto $\mathcal{M}_{\varepsilon,q}$ (using ${\mathsf{KL}}$-divergence), and then sample from the projection. Specifically, their sampling distribution $\hat{p}$ is obtained by solving
@@ 3 Main Results: Minimax Optimal Mechanisms
${\Gamma_{f}(q,\varepsilon)}\coloneqq\inf_{{\mathsf{K}}\in{\mathcal{Q}}_{\varepsilon}\atop q{\mathsf{K}}=q}\sup_{p\in\Delta({\mathcal{X}})}{D_{f}}(p\|p{\mathsf{K}}).$ (4)
The mechanism that solves the above optimization is the minimax optimal mechanism with respect to the choice of $f$-divergence. In the case of ${\mathsf{TV}}$-distance, such optimal mechanism has a practical interpretation. Let $X\sim p$ and $Z$ is the output of an $\varepsilon$-locally private sampler. Optimal mechanism attaining $\Gamma_{{\mathsf{TV}}}(q,\varepsilon)$ corresponds to the best sampler that minimizes the worst-case $\Pr(X\neq Z)$.
@@ 3.1 Optimal Utility
${\Gamma_{f}(q,\varepsilon)}$ $=\frac{1-q_{\min}}{e^{\varepsilon}q_{\min}+1-q_{\min}}f(0)$ (5)
$+\frac{e^{\varepsilon}q_{\min}}{e^{\varepsilon}q_{\min}+1-q_{\min}}f\left(\frac{e^{\varepsilon}q_{\min}+1-q_{\min}}{e^{\varepsilon}q_{\min}}\right).$
@@ 3.2 Optimal Mechanisms
$\sup_{p\in\Delta({\mathcal{X}})}{D_{f}}(p\|p{\mathsf{K}})$ $=\max_{i}{D_{f}}(\delta_{i}\|\delta_{i}{\mathsf{K}})$ (9)
$=f(0)(1-{\mathsf{K}}_{\min})+{\mathsf{K}}_{\min}f(\frac{1}{{\mathsf{K}}_{\min}}).$
@@
Figure 1: Visualization of a single recursion step in Algorithm 1.
@@
$\sup_{p\in\Delta({\mathcal{X}})}{D_{f}}(p\|p{\mathsf{K}}_{q,\varepsilon})=\frac{1-q_{\min}}{e^{\varepsilon}q_{\min}+1-q_{\min}}f(0)$ (14)
$\quad+\frac{e^{\varepsilon}q_{\min}}{e^{\varepsilon}q_{\min}+1-q_{\min}}f\left(\frac{e^{\varepsilon}q_{\min}+1-q_{\min}}{e^{\varepsilon}q_{\min}}\right).$
@@
It is important to note that our algorithm achieves a time complexity of $O(n^{3})$. This is a significant improvement over the naive linear programming approach for solving (4) in the case of ${\mathsf{TV}}$-distance. The naive method has a complexity of $O\left(n^{5}\log\left(\frac{n}{\delta}\right)\right)$ (Cohen et al., 2021), where $n$ is the support size of $p$ and $q$, and $\delta$ is the desired accuracy. With an $O(n^{3})$ complexity, our method scales efficiently to much larger datasets.
@@
Let ${\mathcal{X}}$ be a finite set and let $q_{\text{u}}$ represent the uniform distribution on ${\mathcal{X}}$. Then, the $n$-ary randomized response mechanism, ${\mathsf{K}}_{\text{RR}}$ (Warner, 1965), is optimal in the sense:
@@
This result confirms that ${\mathsf{K}}_{\text{RR}}$ achieves the optimal utility when the prior is uniform and recovers the utility established by Park et al. (2024), where no public prior was assumed.
@@ 4 Overview of Proofs
${\Gamma_{f}(q,\varepsilon)}\geq\frac{1-q_{\min}}{e^{\varepsilon}q_{\min}+1-q_{\min}}f(0)$ (16)
$+\frac{e^{\varepsilon}q_{\min}}{e^{\varepsilon}q_{\min}+1-q_{\min}}f(\frac{e^{\varepsilon}q_{\min}+1-q_{\min}}{e^{\varepsilon}q_{\min}}).$
@@
${\mathsf{K}}_{\min}f(\frac{1}{{\mathsf{K}}_{\min}})+f(0)(1-{\mathsf{K}}_{\min})$ (18)
$\geq\frac{1-q_{\min}}{e^{\varepsilon}q_{\min}+1-q_{\min}}f(0)$
$+\frac{e^{\varepsilon}q_{\min}}{e^{\varepsilon}q_{\min}+1-q_{\min}}f(\frac{e^{\varepsilon}q_{\min}+1-q_{\min}}{e^{\varepsilon}q_{\min}}).$
@@
$\begin{split}&\sup_{p\in\Delta({\mathcal{X}})}{D_{f}}(p\|p{\mathsf{K}}_{q,\varepsilon})=\frac{1-q_{\min}}{e^{\varepsilon}q_{\min}+1-q_{\min}}f(0)\\ &~{}~{}~{}~{}+\frac{e^{\varepsilon}q_{\min}}{e^{\varepsilon}q_{\min}+1-q_{\min}}f(\frac{e^{\varepsilon}q_{\min}+1-q_{\min}}{e^{\varepsilon}q_{\min}}).\end{split}$ (19)
@@ 5 Experiments
In this section, we compare the performance of our minimax optimal private sampler to the state-of-the-art private sampler, the relative mollifier framework (Definition 2). The evaluation is done on synthetic data and two real-world datasets: Avezu CTR (Kaggle, 2015) and MovieLens 1M (Harper and Konstan, 2015). The code for experiments is available on Github 111 https://github.com/bzamanlooy/LDP-sampling-public-data, and details on the computing resources are given in Section A.11.
@@ 5.1 Computation of Samplers
As for the relative mollifier sampling framework (Definition 2), we interpret the reference distribution in Definition 2 as the public prior. In their framework, we determine the normalizing constant $C$ described in (3) using a bisection approach (Brent, 2013). While Husain et al. (2020) focus exclusively on ${\mathsf{KL}}$-divergence in the relative mollifier framework, we extend their work to include the ${\mathsf{TV}}$-distance, offering an additional benchmark. This extension uses linear programming, with details provided in Appendix A.8.